# The Equational Theories Project: Advancing Collaborative Mathematical Research at Scale

### v1

@@ 7. Proof automation
In this project we used proof automation in two ways: automated theorem provers (ATPs) and Lean tactics. ATPs are generally stand-alone tools that implement a (semi-)decision procedure for a given formal language or related set of languages. For example, Vampire [DBLP:conf/cav/KovacsV13, TheVampireDiary] is an ATP focused primarily on first-order logic using superposition, which we used extensively in this project. We also made extensive use of Prover9 and Mace4 [prover9-mace4].
@@
Tactics in Lean, on the other hand, are meta-programs [DBLP:journals/pacmpl/EbnerURAM17] that build proofs. In other words, they essentially take Lean code as input and produce Lean code as output. In this manner, they look like another keyword in the language, and are tightly integrated by producing proofs directly. Under the hood, their implementation can be arbitrarily complex, from syntactic sugar to full decision procedures. The duper tactic [DBLP:conf/itp/CluneQBA24], for example, implements a superposition calculus similar to Vampire’s, but for dependent types — Lean’s underlying logical foundation.
@@ 7.1. Proof techniques
The two main families of ATPs and tactics we used are based on superposition/saturation and equational reasoning. In this context we also include SMT solvers, which combine specific decision procedures for theories, like congruence closure for equational reasoning, with satisfiability (SAT) solving [deMoura-Bjorner-2009]. Finally, we also used aesop [DBLP:conf/cpp/LimpergF23], which implements a version of tableau search. This was used mainly to help specific constructions in refutations, and is not specific to proving or disproving magma implications in this sense. We describe our use of aesop in Section˜7.2 below.
Saturation. Most of the ATPs used extensively in this project rely primarily on saturation procedures in the superposition calculus. For example, this is the case for Vampire [DBLP:conf/cav/KovacsV13]. [^12] See also [DBLP:journals/cacm/BentkampBNTVW23] for a gentler exposition. The core idea of these provers is that they take a set of assumptions and a conjecture, expressed in (say) first-order logic. The conjecture is negated and added to the set of assumptions, which are all put into a normal form. The ATP then tries to refute the negation by applying rules of an underlying calculus, until a proof of false (a contradiction) is derived. In this case, the conjecture was (classically) true, and the ATP has found a proof by contradiction, often called a “refutation” or “saturation” proof.
@@
In this project we used Vampire [DBLP:conf/cav/KovacsV13], Duper [DBLP:conf/itp/CluneQBA24] and Prover9 and Mace4 [prover9-mace4] which are all based on variants of saturation for proving.
Equational Reasoning. As already discussed in Section˜6.3, equational reasoning is a type of reasoning that is based [^13] More precisely, one can formalize this reasoning using Birkhoff’s five rules of inference (reflexive, symmetric, transitive, replacement, and substitution); see, e.g., [burris]. on equational logic and rewriting with congruence [term-rewriting]. In general, an equational reasoning procedure takes a series of equations and tries to determine whether another equation can be deduced from it. A core tool in equational reasoning are e-graphs, a data structure used to represent congruence classes of terms. By themselves, e-graphs provide an efficient means of implementing a decision procedure for congruence closure over ground equations (i.e., equations without variables). Extensions to this procedure, for example by quantifier instantiation via e-matching [DBLP:conf/cade/MouraB07], also allow for a semi-decision procedure for congruence closure over non-ground equations.
SMT solvers like Z3 [DBLP:conf/tacas/MouraB08] use equational reasoning for deciding the theory of equality with uninterpreted functions [DBLP:series/txtcs/KroeningS16, DBLP:conf/cade/MouraB07]. On the other hand, equality saturation [DBLP:journals/pacmpl/WillseyNWFTP21] uses e-graphs by extending congruence closure to a more controlled search, enabling optimization and conditional rewriting. One of the main advantages of equational reasoning for implications of magma laws is that we get very explicit proofs: a proof that $l\models l^{\prime}$ is given by a sequence of rewrites that starts at the left-hand side of $l^{\prime}$ and arrives at the right-hand side through applications of $l$.
In this project we used Z3 [DBLP:conf/tacas/MouraB08], Prover9 and Mace4 [prover9-mace4], a custom ATP MagmaEgg for magmas based on egg [DBLP:journals/pacmpl/WillseyNWFTP21], and the Lean egg tactic [DBLP:journals/pacmpl/KoehlerGBGTS24, rossel2024equality], which all work with equational logic. We have also reasoned with manual (custom written) heuristics about simple rewrites.
@@ 7.2. Integration of automation procedures
Proof Reconstruction. The relative simplicity of the objects used in this project benefits the implementations of proof reconstruction. By focusing on the given problem domain, difficult reconstruction issues, like complex dependent types, could be ignored.
@@
For equational proofs from external provers, like MagmaEgg, we also used a tailored version of reconstruction. Specifically, the MagmaEgg implementation turns explanations [nieuwenhuis2005proof] from egg into Lean proofs by simple applications of the defining properties of equality as shown in Figure˜11.
@@
Semi-Automated Counterexample Guidance. Another use of ATPs has been in a semi-automatic fashion, to find counterexamples. The general strategy was to use ATPs to find counterexamples to implications by building magmas iteratively. If we want to build a counterexample to $l\models l^{\prime}$, we want to construct a magma where $l$ holds but $l^{\prime}$ does not. In this method, we iteratively strengthen a construction with additional hypotheses, and use the ATP to check whether these hypotheses are not too strong (to imply $l^{\prime}$) or unsound (to disallow $l$).
While equational reasoning can also be used in a semi-automatic fashion to prove equations [DBLP:journals/pacmpl/KoehlerGBGTS24], the positive implications in the main implication graph of the project were all simple enough that we did not need a semi-automatic approach for them.
@@ 7.3.1. Employing several ATPs
In general, given a batch of problems of interest, it is useful to employ several ATPs when they complement each other on the batch — i.e., when they can solve different sets of problems. This is often the case with Vampire and Prover9: e.g., in a 2024 study [LPAR2024:Prover9_Unleashed_Automated_Configuration], it is shown that from a batch of around 770 TPTP problems solved with Vampire and Prover9 together (with some restrictions), around $60\%$ are solved by both, $20\%$ are solved only by Vampire, and the remaining $20\%$ are solved only by Prover9.171717In contrast, in said study the E prover solves only a single problem not solved by Vampire.
@@ 7.3.3. Flow control
Vampire is equipped with a user-friendly and powerful standard mode, the CASC mode (or CASC SAT mode for finite model building) that in many situations avoids the need of configuring specific search options for the given problem. This mode invokes a sequence of strategies (different option configurations) with assigned time limits. The time given to each strategy is important: counterintuitively, giving less time to the CASC mode may end up producing a faster proof: there are proofs that can be found when the time limit is set to less than $5\%$ of the time Vampire needs to find a proof without the time limit [DBLP:conf/cav/KovacsV13]. At least two factors contribute to this effect: the correct strategy is explored sooner because less time is spent on each strategy, and the behaviour of Vampire’s default saturation algorithm, the limited resource strategy (LRS) algorithm.
@@ 7.3.5. The weight-sos limit strategy
Vampire does not allow the imposition of a user-defined maximum weight, although its default LRS algorithm already implements a dynamic weight limit. According to [janota2025experimentalresultsvampireequational], Vampire 5.0 can prove all true implications from the ETP, and can prove 99.97% of them using less than 500 instructions per proof.
@@ 7.3.6. Mace4
Configuration: (2,4,Y/N) (0,4,Y/N) (2,3,Y/N) (2,2,Y/N) (2,1,Y/N)
Size: 14 (in 127s) 14 (in 166s) 13 (in 290s) 9 (in 33s) 7 (in 8s)
@@
Table 3. Time to find one model of $\mathrm{E}677$ of size $9$ and to exhaust the same size for different configurations. Some finite-setting formulas were included to speed up the search (see Section˜7.3.8). The configurations not listed here were unable to find a model in 4000s.
Configuration One model Exhaust size
(2,1,N) 16s 190s
(2,1,Y) 16s 210s
(0,1,Y) 20s 149s
(0,1,N) 32s 169s
(2,2,Y) 154s 72000s+
(2,2,N) 180s 43000s+
(2,3,N) 557s 877s
(2,3,Y) 587s 941s
(2,4,Y) 596s 761s
(2,4,N) 620s 730s
(0,3,N) 3746s 4151s
(0,3,Y) 3777s 4254s
@@
Table 4. Comparison of the two best configurations for finding certain special $\mathrm{E}677$ models, both for exhausting each size and for finding a single model, with best times in boldface. Note that initially (2,4,N) seems the best configuration, but in the long run, (2,3,Y) outperforms it. The comparison is particularly misleading at size 16 when searching for a single model.
Size (2,3,Y) exhaust (2,4,N) exhaust (2,3,Y) model (2,4,N) model
9 0.14s 0.04s 0.02s 0.0s
10 0.32s 0.12s - -
11 0.82s 0.37s 0.15s 0.01s
12 2.17s 1.2s - -
13 6.18s 4.35s 2.74s 0.91s
14 15.47s 14.39s - -
15 38.82s 50.87s - -
16 120s 175s 11.95s 0.02s
17 306s 660s - -
18 883s 2382s - -
19 2300s 3600s+ 846s 2082s
@@ 9. Spectrum of equational laws
Several extensions will be described in a separate publication: determining the spectrum and not only whether it is full; the spectrum of simple magmas or (sub)directly irreducible magmas; tracking multiplicity, namely counting (or bounding asymptotically) how many finite magmas exist of each size in the spectrum; this is referred to as the fine spectrum in [taylor]. These detailed considerations reveal profound differences in how much an equational law constrains the magma operation, and may help organize the implication graph into different families.
@@
  • For $(a,b)=(0,0)$ the operation is the constant operation, which is a model of any law whose sides both have positive order. Equivalently, these laws are consequences of the constant law $\mathrm{E}41$.
@@
  • The operation with $(a,b)=(1,1)$ cannot satisfy a law for all $n$.
@@ 10.1. Describing groups as magmas
The ETP is focused exclusively on magmas, which only feature a single (binary) operation. Many mathematical structures traditionally defined using several operations can nevertheless be fully described as magmas with a well-chosen combined operation, from which the whole structure can be reconstructed. The first example is how Boolean algebras defined in terms of three operations $(\land,\lor,\lnot)$ were equivalently described in 1913 in terms of the Sheffer stroke $x\diamond y\coloneqq\lnot(x\land y)$ [sheffer]. Once such a single operation is found, a separate endeavor is to determine which laws it must satisfy to get the desired structure, and, in favorable cases find a single law that encapsulates the whole structure, or even find all equivalent laws of minimum order. The earliest such example is Tarski’s description of abelian groups in terms of subtraction $x\diamond y\coloneqq x+(-y)$, subject to a single axiom $\mathrm{x}\simeq\mathrm{y}\diamond(\mathrm{z}\diamond(\mathrm{x}\diamond(\mathrm{y}\diamond\mathrm{z})))$ (E543), found in 1938 [Tarski1938]. It then took three decades [higman-neumann, Sholander01021959, Padmanabhan_1969] to sort out the full equivalence class of $\mathrm{E}543$ among laws of order $4$. For Boolean algebras, a minimum-order single-law description was only found in [mccune_et_al], nine decades after Sheffer’s work.
@@
Here, we describe the case of groups. The binary operation *, unary operation $(\cdot)^{-1}$, and nullary operation $e$ (identity element) can be repackaged into a single division operation $x\diamond y\coloneqq x*y^{-1}$, from which the original operations are easily reconstructed: for instance $x*y=x\diamond((y\diamond y)\diamond y)$. A group equipped with division, called a Ward quasigroup, is a magma $(G,\diamond)$ satisfying the unipotence law $\mathrm{x}\diamond\mathrm{x}\simeq\mathrm{y}\diamond\mathrm{y}$ (E40), the right-unit squares law $\mathrm{x}\simeq\mathrm{x}\diamond(\mathrm{y}\diamond\mathrm{y})$ (E11), and a version of the associativity law dubbed the half-group law, $\mathrm{x}\diamond\mathrm{y}\simeq(\mathrm{x}\diamond\mathrm{z})\diamond(\mathrm{y}\diamond\mathrm{z})$ (E3737), from which group axioms are easily derived. These three laws are equivalent to a single law $\mathrm{E}_{\mathrm{HN}}\coloneqq\mathrm{E}42323216$ of order $8$, found by Higman and Neumann [higman-neumann],
@@
McCune found two more laws equivalent to this one and of the same order [mccune1993single], $\mathrm{E}42302852$ and $\mathrm{E}147976245$. A natural question is to find all characterizations of Ward quasigroups (groups equipped with division) with minimum order. Throughout our exploration, we used two criteria: the law must be satisfied by group division, and must fail for magmas that are not Ward quasigroups.

### v2

@@ 7. Proof automation
In this project we used proof automation in two ways: automated theorem provers (ATPs) and Lean tactics. ATPs are generally stand-alone tools that implement a (semi-)decision procedure for a given formal language or related set of languages. For example, Vampire [38, 8] is an ATP focused primarily on first-order logic using superposition, which we used extensively in this project. We also made extensive use of Prover9 and Mace4 [43].
@@
Tactics in Lean, on the other hand, are meta-programs [24] that build proofs. In other words, they essentially take Lean code as input and produce Lean code as output. In this manner, they look like another keyword in the language, and are tightly integrated by producing proofs directly. Under the hood, their implementation can be arbitrarily complex, from syntactic sugar to full decision procedures. The duper tactic [17], for example, implements a superposition calculus similar to Vampire’s, but for dependent types — Lean’s underlying logical foundation.
@@ 7.1. Proof techniques
The two main families of ATPs and tactics we used are based on superposition/saturation and equational reasoning. In this context we also include SMT solvers, which combine specific decision procedures for theories, like congruence closure for equational reasoning, with satisfiability (SAT) solving [19]. Finally, we also used aesop [41], which implements a version of tableau search. This was used mainly to help specific constructions in refutations, and is not specific to proving or disproving magma implications in this sense. We describe our use of aesop in Section˜7.2 below.
7.1.1. Saturation
Most of the ATPs used extensively in this project rely primarily on saturation procedures in the superposition calculus. For example, this is the case for Vampire [38]. [^10] See also [9] for a gentler exposition. The core idea of these provers is that they take a set of assumptions and a conjecture, expressed in (say) first-order logic. The conjecture is negated and added to the set of assumptions, which are all put into a normal form. The ATP then tries to refute the negation by applying rules of an underlying calculus, until a proof of false (a contradiction) is derived. In this case, the conjecture was (classically) true, and the ATP has found a proof by contradiction, often called a “refutation” or “saturation” proof.
@@ 7.1.1. Saturation
In this project we used Vampire [38], Duper [17] and Prover9 and Mace4 [43] which are all based on variants of saturation for proving.
7.1.2. Equational Reasoning
As already discussed in Section˜6.3, equational reasoning is a type of reasoning that is based [^11] More precisely, one can formalize this reasoning using Birkhoff’s five rules of inference (reflexive, symmetric, transitive, replacement, and substitution); see, e.g., [14]. on equational logic and rewriting with congruence [7]. In general, an equational reasoning procedure takes a series of equations and tries to determine whether another equation can be deduced from it. A core tool in equational reasoning are e-graphs, a data structure used to represent congruence classes of terms. By themselves, e-graphs provide an efficient means of implementing a decision procedure for congruence closure over ground equations (i.e., equations without variables). Extensions to this procedure, for example by quantifier instantiation via e-matching [20], also allow for a semi-decision procedure for congruence closure over non-ground equations.
SMT solvers like Z3 [21] use equational reasoning for deciding the theory of equality with uninterpreted functions [39, 20]. On the other hand, equality saturation [69] uses e-graphs by extending congruence closure to a more controlled search, enabling optimization and conditional rewriting. One of the main advantages of equational reasoning for implications of magma laws is that we get very explicit proofs: a proof that $l\models l^{\prime}$ is given by a sequence of rewrites that starts at the left-hand side of $l^{\prime}$ and arrives at the right-hand side through applications of $l$.
In this project we used Z3 [21], Prover9 and Mace4 [43], a custom ATP MagmaEgg for magmas based on egg [69], and the Lean egg tactic [37, 60], which all work with equational logic. We have also reasoned with manual (custom written) heuristics about simple rewrites.
@@ 7.2. Integration of automation procedures
7.2.1. Proof Reconstruction
The relative simplicity of the objects used in this project benefits the implementations of proof reconstruction. By focusing on the given problem domain, difficult reconstruction issues, like complex dependent types, could be ignored.
@@ 7.2.1. Proof Reconstruction
@[equational_result] theorem Equation999_implies_Equation86
(G : Type*) [Magma G] (h : Equation999 G) : Equation86 G := by
by_contra nh
simp only [not_forall] at nh
[…]
have eq9 (X0 X1 X2 X3 : G) : (X1 $\diamond$ ((X2 $\diamond$ X3) $\diamond$ (X1 $\diamond$ X0))) = X0 :=
mod_symm (h ..)
have eq10 : sK0 $\neq$ (sK1 $\diamond$ (sK2 $\diamond$ (sK1 $\diamond$ sK0))) := mod_symm nh
have eq11 (X3 X4 X5 : G) : (X4 $\diamond$ (X3 $\diamond$ (X4 $\diamond$ X5))) = X5 := superpose eq9 eq9
have eq21 : sK0 $\neq$ sK0 := superpose eq11 eq10
subsumption eq21 rfl
@@
For equational proofs from external provers, like MagmaEgg, we also used a tailored version of reconstruction. Specifically, the MagmaEgg implementation turns explanations [56] from egg into Lean proofs by simple applications of the defining properties of equality as shown in Figure˜11.
@[equational_result] theorem Equation3973_implies_Equation4023
(G : Type _) [Magma G] (h : Equation3973 G) : Equation4023 G := fun x y z =>
let v0 := M z x
let v1 := M z v0
let v2 := M v1 y
let v3 := M z v1
have h4 := R v2
have h5 := R z
have h6 := h z x v1
let v7 := M x (M v1 z)
T (T (T (T (h x y z) (h (M y v0) z v2)) (C (C h5 (C h4 (T (T (h y v0 v2)
(C (C (T (T (T h6 (h v7 v1 v0)) (C (C (R v1) (T (h v0 v7 z) (C (S h6) h5)))
(R v0))) (S (h z v1 v0))) (R (M v2 y))) h4)) (S (h y v3 v2))))) h4))
(S (h (M y v3) z v2))) (S (h v1 y z))
@@
7.2.2. Semi-Automated Counterexample Guidance
Another use of ATPs has been in a semi-automatic fashion, to find counterexamples. The general strategy was to use ATPs to find counterexamples to implications by building magmas iteratively. If we want to build a counterexample to $l\models l^{\prime}$, we want to construct a magma where $l$ holds but $l^{\prime}$ does not. In this method, we iteratively strengthen a construction with additional hypotheses, and use the ATP to check whether these hypotheses are not too strong (to imply $l^{\prime}$) or unsound (to disallow $l$).
While equational reasoning can also be used in a semi-automatic fashion to prove equations [37], the positive implications in the main implication graph of the project were all simple enough that we did not need a semi-automatic approach for them.
@@ 7.3.1. Employing several ATPs
In general, given a batch of problems of interest, it is useful to employ several ATPs when they complement each other on the batch — i.e., when they can solve different sets of problems. This is often the case with Vampire and Prover9: e.g., in a 2024 study [2], it is shown that from a batch of around 770 TPTP problems solved with Vampire and Prover9 together (with some restrictions), around $60\%$ are solved by both, $20\%$ are solved only by Vampire, and the remaining $20\%$ are solved only by Prover9.151515In contrast, in said study the E prover solves only a single problem not solved by Vampire.
@@ 7.3.3. Flow control
Vampire is equipped with a user-friendly and powerful standard mode, the CASC mode (or CASC SAT mode for finite model building) that in many situations avoids the need of configuring specific search options for the given problem. This mode invokes a sequence of strategies (different option configurations) with assigned time limits. The time given to each strategy is important: counterintuitively, giving less time to the CASC mode may end up producing a faster proof: there are proofs that can be found when the time limit is set to less than $5\%$ of the time Vampire needs to find a proof without the time limit [38]. At least two factors contribute to this effect: the correct strategy is explored sooner because less time is spent on each strategy, and the behaviour of Vampire’s default saturation algorithm, the limited resource strategy (LRS) algorithm.
@@ 7.3.5. The weight-sos limit strategy
Vampire does not allow the imposition of a user-defined maximum weight, although its default LRS algorithm already implements a dynamic weight limit. According to [32], Vampire 5.0 can prove all true implications from the ETP, and can prove 99.97% of them using less than 500 instructions per proof.
@@ 7.3.6. Mace4
Configuration (2,4,Y/N) (0,4,Y/N) (2,3,Y/N) (2,2,Y/N) (2,1,Y/N)
Size 14 (in $127\text{\,}\mathrm{s}$) 14 (in $166\text{\,}\mathrm{s}$) 13 (in $290\text{\,}\mathrm{s}$) 9 (in 33 $3\text{\,}\mathrm{s}$) 7 (in $8\text{\,}\mathrm{s}$)
@@
Table 3. Time to find one model of $\mathrm{E}677$ of size $9$ and to exhaust the same size for different configurations. Some finite-setting formulas were included to speed up the search (see Section˜7.3.8). The configurations not listed here were unable to find a model in $4000\text{\,}\mathrm{s}$.
Configuration One model Exhaust size
(2,1,N) $16\units$ $190\units$
(2,1,Y) $16\units$ $210\units$
(0,1,Y) $20\units$ $149\units$
(0,1,N) $32\units$ $169\units$
(2,2,Y) $154\units$ $72\,000\units$
(2,2,N) $180\units$ $43\,000\units$
(2,3,N) $557\units$ $877\units$
(2,3,Y) $587\units$ $941\units$
(2,4,Y) $596\units$ $761\units$
(2,4,N) $620\units$ $730\units$
(0,3,N) $3746\units$ $4151\units$
(0,3,Y) $3777\units$ $4254\units$
Table 4. Comparison of the two best configurations for finding certain special $\mathrm{E}677$ models, both for exhausting each size and for finding a single model, with best times in boldface. Note that initially (2,4,N) seems the best configuration, but in the long run, (2,3,Y) outperforms it. The comparison is particularly misleading at size 16 when searching for a single model. All times are in seconds.
Exhaustion Finding one model
Size (2,3,Y) (2,4,N) (2,3,Y) (2,4,N)
9 $0.14$ $0.04$ $0.02$ $0.0$
10 $0.32$ $0.12$
11 $0.82$ $0.37$ $0.15$ $0.01$
12 $2.17$ $1.2$
13 $6.18$ $4.35$ $2.74$ $0.91$
14 $15.47$ $14.39$
15 $38.82$ $50.87$
16 $120$ $175$ $11.95$ $0.02$
17 $306$ $660$
18 $883$ $2382$
19 $2300$ >>$3600$ $846$ $2082$
@@ 9. Spectrum of equational laws
Several extensions will be described in a separate publication: determining the spectrum and not only whether it is full; the spectrum of simple magmas or (sub)directly irreducible magmas; tracking multiplicity, namely counting (or bounding asymptotically) how many finite magmas exist of each size in the spectrum; this is referred to as the fine spectrum in [67]. These detailed considerations reveal profound differences in how much an equational law constrains the magma operation, and may help organize the implication graph into different families.
@@
  • For $(a,b)=(0,0)$ the operation is the constant operation, which is a model of $\mathrm{E}1$ and any law whose sides both have positive order. Equivalently, these laws are consequences of the constant law $\mathrm{E}46$.
@@
  • The operation for $(a,b)=(1,1)$ is abelian group addition, which is useless as any law it satisfies for all $n$ is also satisfied by the constant operation $(a,b)=(0,0)$.
@@ 10.1. Describing groups as magmas
The ETP is focused exclusively on magmas, which only feature a single (binary) operation. Many mathematical structures traditionally defined using several operations can nevertheless be fully described as magmas with a well-chosen combined operation, from which the whole structure can be reconstructed. The first example is how Boolean algebras defined in terms of three operations $(\land,\lor,\lnot)$ were equivalently described in 1913 in terms of the Sheffer stroke $x\diamond y\coloneqq\lnot(x\land y)$ [61]. Once such a single operation is found, a separate endeavor is to determine which laws it must satisfy to get the desired structure, and, in favorable cases find a single law that encapsulates the whole structure, or even find all equivalent laws of minimum order. The earliest such example is Tarski’s description of abelian groups in terms of subtraction $x\diamond y\coloneqq x+(-y)$, subject to a single axiom $\mathrm{x}\simeq\mathrm{y}\diamond(\mathrm{z}\diamond(\mathrm{x}\diamond(\mathrm{y}\diamond\mathrm{z})))$ (E543), found in 1938 [66]. It then took three decades [29, 62, 57] to sort out the full equivalence class of $\mathrm{E}543$ among laws of order $4$. For Boolean algebras, a minimum-order single-law description was only found in [47], nine decades after Sheffer’s work.
@@
Here, we describe the case of groups. The binary operation *, unary operation $(\cdot)^{-1}$, and nullary operation $e$ (identity element) can be repackaged into a single division operation $x\diamond y\coloneqq x*y^{-1}$, from which the original operations are easily reconstructed: for instance $x*y=x\diamond((y\diamond y)\diamond y)$. A group equipped with division, called a Ward quasigroup, is a magma $(G,\diamond)$ satisfying the unipotence law $\mathrm{x}\diamond\mathrm{x}\simeq\mathrm{y}\diamond\mathrm{y}$ (E40), the right-unit squares law $\mathrm{x}\simeq\mathrm{x}\diamond(\mathrm{y}\diamond\mathrm{y})$ (E11), and a version of the associativity law dubbed the half-group law, $\mathrm{x}\diamond\mathrm{y}\simeq(\mathrm{x}\diamond\mathrm{z})\diamond(\mathrm{y}\diamond\mathrm{z})$ (E3737), from which group axioms are easily derived. These three laws are equivalent to a single law $\mathrm{E}_{\mathrm{HN}}\coloneqq\mathrm{E}42323216$ of order $8$, found by Higman and Neumann [29],
@@
McCune found two more laws equivalent to this one and of the same order [48], $\mathrm{E}42302852$ and $\mathrm{E}147976245$. A natural question is to find all characterizations of Ward quasigroups (groups equipped with division) with minimum order. Throughout our exploration, we used two criteria: the law must be satisfied by group division, and must fail for magmas that are not Ward quasigroups.